OpenStudy (anonymous):

What is the simplified form of (x^2-5x+6/15xy^2) / (2x^2-7x+3/5x^2y)?

OpenStudy (campbell_st):

factor the quadratics, then apply the flip and multiply rule for dividing by fractions. so you get \[\frac{(x -2)(x-3)}{15xy^2} \times \frac{5x^2y}{(2x -1)(x - 3)}\] now look for common factors that can be removed.

OpenStudy (anonymous):

x-3 you can cancel

OpenStudy (campbell_st):

yes thats one... and you'll find a couple of other thinks... common factors of the denominator in the 1st fraction and numerator in the 2nd fraction

OpenStudy (anonymous):

cant you also take out a 5 an x and a y ?

OpenStudy (campbell_st):

thats it... then multiply what you have left for the solution

OpenStudy (anonymous):

x^2-2/6xy-3y?

OpenStudy (campbell_st):

not quite \[\frac{x(x -2)}{3y(2x - 1)}\] then distribute, your error is in the numerator

OpenStudy (anonymous):

or just keep it simplelike x(x-2)/3y(2x-1) yeah thats what i got x^2-2/6xy-3y

OpenStudy (anonymous):

oh -2x

OpenStudy (campbell_st):

yep... thats it

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (campbell_st):

just close this and post another, there are plenty of people to help.,